OpenStudy (anonymous):

Calculate the following integrals \[\Large \int_0^a dx x^{2}e^{x}\]

sam (.sam.):

Integration by parts 2 times

OpenStudy (anonymous):

Sam, a complete solution would be much apreciate, i have exam tomorrow and its possibly be on my exam or very smilar question of this..

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+x^2+e^x hit show steps

sam (.sam.):

let u=x^2 dv=e^x du=2xdx v=e^x \[x^2e^x-\int\limits_{}^{}e^x(2x)dx\] let u=2x dv=e^x du=2dx v=e^x \[x^2e^x-(2xe^x-\int\limits_{}^{}2e^xdx)\] \[x^2e^x-(2xe^x-2e^x)\]

OpenStudy (anonymous):

thank you Mr Saab and thank you Sam

sam (.sam.):

yw